# Low-Dimensional Adaptation of Rectified Flow: A New Perspective through the Lens of Diffusion and Stochastic Localization

### v1

@@ 4.1 Low-dimensionality of target distribution
To formalize the notion of low-dimensionality of the target data generating distribution $p_{1}$, we turn to the geometric complexity of $\mathcal{X}:={\sf supp}(p_{1})\subseteq\mathbb{R}^{d}$ as measured by its metric entropy.
@@
Let $\epsilon_{0}=N^{-c_{\epsilon_{0}}}$ for some sufficiently large $c_{\epsilon_{0}}>0$. The $\epsilon_{0}$ covering number of $\mathcal{X}$ with respect to the Euclidean norm, $\mathcal{N}(\mathcal{X},\|\cdot\|_{2},\epsilon_{0})$, satisfies
$\log\mathcal{N}(\mathcal{X},\|\cdot\|_{2},\epsilon_{0})\leq C_{\rm cover}k\log N,$
for some constant $C_{\rm cover}>0$ and integer $k$. The quantity $k$ is referred as the intrinsic-dimension of $\mathcal{X}$.
@@
There exists a universal constant $c_{R}>0$ such that
$\sup_{x\in\mathcal{X}}\|x\|_{2}\leq R,\text{where $R=N^{c_{R}}$}.$
The bounded support assumption is not as stringet as it may appear at first, as the diameter of the support $\mathcal{X}$ is allowed to scale polynomially (with arbitrarily large degree) in the number of iterations of the sampler, which accommodates a very wide range of practical applications, such as image generation.
@@ 4.2 Assumptions on velocity approximation error
Our next set of assumptions are concerned with the approximation error of the estimated velocity field $\hat{v}_{t}$ and are also standard.
@@
Define $(\varepsilon^{\rm v}_{i})^{2}:=\mathbb{E}\|v_{t_{i}}(X_{t_{i}})-\widehat{v}_{t_{i}}(X_{t_{i}})\|_{2}^{2}$. Then,
$\frac{1}{N}\sum_{i=1}^{N}(\varepsilon^{\rm v}_{i})^{2}\leq\varepsilon_{\rm v}^{2}.$
The above assumption essentially requires a good control on the mean-squared error of the estimated velocity field. Similar assumptions have been considered in (Bansal et al., 2025; Guan et al., 2025) to prove Wasserstein convergence of RF. In the context of score-matching algorithms, similar assumptions related to the score-approximation error have also appeared in prior works such as (Benton et al., 2023; Chen et al., 2023; Li et al., 2024b; Liang et al., 2025).
Finally, to derive convergence rates of the deterministic samplers of RF that are adaptive to the intrinsic dimension of the support of the target distribution, we require the velocity estimator $\widehat{v}_{t}$ to have higher order accuracy guarantees.
@@
    Let $(\varepsilon^{\rm J,1}_{i})^{2}:=\mathbb{E}\|\nabla v_{t_{i}}(X_{t_{i}})-\nabla\widehat{v}_{t_{i}}(X_{t_{i}})\|_{F}^{2}$. Then, $N^{-1}\sum_{i=1}^{N}(\varepsilon^{\rm J,1}_{i})^{2}\leq\varepsilon_{\rm J,1}^{2}$. Also, we assume
    $\sup_{x\in\mathbb{R}^{d}}\eta_{i}\|\nabla v_{t_{i}}(x)-\nabla\widehat{v}_{t_{i}}(x)\|_{{\rm op}}\leq 1/8$
    for all $0\leq i\leq N-2$.
@@
Nearly identical assumptions have been used in Liang et al. (2025) in the context of DDIM. Intuitively, an adequate control of the higher-order approximation terms mitigates the propagation of error during sampling, as the deterministic samplers are unable to self-correct the path of the flow. In contrast, Assumption 4.4 is not required to prove convergence guarantees for Stoc-RF sampler as the Langevin term automatically acts as a path corrector term.
@@ 4.3 Convergence rate of rectified flow
In this section, we will establish the main convergence theorem for RF sampler (5), $Y_{t_{i+1}}=Y_{t_{i}}+\eta_{i}\widehat{v}_{t_{i}}(Y_{t_{i}}),\;Y_{0}\sim N(0,I_{d}).$
@@
Our convergence guarantees for RF depend crucially on a carefully constructed U-shaped time schedule $\{t_{i}\}_{i=0}^{N}$, which we describe next. For given, user-specified values $h\in(0,1)$ and $\delta\geq 0$, we consider the non-uniform time partition
@@
We remark that the proposed time schedule is practicable, as it does not depend on any properties of $p_{1}$. Typically, $\delta$ is chosen to be small, as this give rise to a U-shaped time discretization as shown in Figure 2. This type of choice is not unusual, as it yields more accurate estimates $\widehat{v}_{t}$ during training when $t\approx 0$ and $t\approx 1$. Prior empirical work by Lee et al. (2024) has shown that such U-shaped choice of partition results in better training of the RF model. Non-uniform time discretizations have also been considered by Huang et al. (2024); Li and Yan (2024); Liang et al. (2025) in the context DDPM and DDIM to establish low-dimensional adaptation of the models. It is important to emphasize that the specific peculiar form of the time schedule (15), while conforming to best empirical practices, is theoretically grounded, and it was derived directly as a result of our technical analysis of the RF dynamics (see Step 7 of the proof of Theorem 4.5, Appendix B.2). We are now ready to present the main convergence result for the RF sampler. The detailed proof of the result is deferred to Appendix B.2.
@@
The dependence on $\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}$. The upper bound in Theorem 4.5 contains a term involving $\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}$, which even for data supported on manifolds would linearly in the ambient dimension $d$. However, in practice one can choose the free parameter $\delta$ to be e.g. $O(d^{-1})$, so that the term becomes negligible. With this choice of $\delta$, last term in (17) scales as $d^{2}N^{-10}$ which is also negligible in practice. For example, if $d=256\times 256$ and $N=80$, then $d^{2}N^{-10}\approx 4\times 10^{-10}$.
@@
Higher-order approximation error terms. The upper bound in Theorem 4.5 is heavily dependent on the higher-order approximation errors of $\widehat{v}_{t}$. As mentioned before, these terms are required to be small as deterministic samplers are unable to self-correct their trajectories. In addition, at $t=0$, the true drift $v_{0}(x)=\mu_{1}-x$ can be estimated by $\widehat{v}_{0}(x)=n^{-1}\sum_{i\in[n]}{X_{1}^{(i)}}-x$, where $\{X_{1}^{(i)}\}_{i\in[n]}\overset{i.i.d.}{\sim}p_{1}$. With this choice of $\widehat{v}_{0}$, we have $\varepsilon^{\rm J,1}_{0}=\varepsilon^{\rm J,2}_{0}=\varepsilon^{\rm H}_{0}=0$, and $\varepsilon^{\rm v}_{0}=\sqrt{\frac{\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}}{n}}$.
@@ 4.4 Convergence rate of Stoc-RF
In this section, we will leverage the equivalence between DDPM and stoc-RF to design an SDE-based sampler that also adapts to the low-dimensionality of $p_{1}$. We begin with the solution $\left\{{Y^{\prime}}\right\}_{\tau\geq 0}$ to the forward process (11). We also consider the time-change map $\tau\in[0,\infty)\mapsto t(\tau)=\frac{\sqrt{\omega_{\tau}}}{\sqrt{\omega_{\tau}}+\sqrt{1-\omega_{\tau}}}.$ In light of (14) and Proposition 3.4, we note that $\left\{{Y^{\prime}_{\tau}}\right\}_{\tau\geq 0}$ and $\{\tilde{Z}_{t}\}_{t\in[0,1]}$ (recall (6)) are equivalent under the time change $\tau\mapsto t(\tau)$, and we have $\frac{Y^{\prime}_{\tau}}{\sqrt{\omega_{\tau}}}\overset{d}{=}\frac{\tilde{Z}_{t(\tau)}}{t(\tau)}$.
In practice, we simulate the forward process by discretizing $\tau$ and setting specific values for $\omega_{\tau}$. In this section, we let $\omega_{\tau}=\prod_{j=1}^{\tau}\alpha_{j}$ with the $\alpha_{j}$’s defined implicitly as
$\beta_{1}=1-\alpha_{1}=\frac{1}{N^{c_{0}}},$ (18)
@@
for a user-specified choice of $c_{0},c_{1}>0$. The time schedule above is very similar to the ones considered by Potaptchik et al. (2024); Li et al. (2024b); Li and Yan (2024); Huang et al. (2024); Liang et al. (2025) to analyze the convergence properties of DDPM. Next, we recall the DDPM sampler considered in Liang et al. (2025); Huang et al. (2024); Li and Yan (2024):
@@
and $\widehat{s}_{Y^{\prime}_{\tau}}$ is an approximate score of $Y^{\prime}_{\tau}$. The provide some intuition behind this sampler, we recall that $1-\alpha_{\tau}=\beta_{\tau}\approx 0$. Hence, $1/\sqrt{\alpha_{\tau}}\approx 1+\beta_{\tau}/2,\delta_{\tau}/\sqrt{\alpha_{\tau}}\approx\beta_{\tau}/2$, and $\nu_{\tau}/\sqrt{\alpha_{\tau}}\approx\sqrt{\beta_{\tau}}$. Substituting this in (19), we get
$\hat{Y}^{\prime}_{\tau-1}\approx\hat{Y}^{\prime}_{\tau}+\frac{\beta_{\tau}}{2}\left\{{\hat{Y}^{\prime}_{\tau}+2\widehat{s}_{Y^{\prime}_{\tau}}(\hat{Y}^{\prime}_{\tau})}\right\}+\sqrt{2.\left({\frac{\beta_{\tau}}{2}}\right)}\xi_{\tau},$ (20)
@@
Therefore, we have $Y^{\prime}_{\tau}\overset{d}{=}\tilde{Z}_{t(\tau)}/\sigma_{t(\tau)}$, and we have $s_{Y^{\prime}_{\tau}}(y)=\sigma_{t(\tau)}s_{t(\tau)}(\sigma_{t(\tau)}y)$. We also define $t_{i}:=t(N-i)$ for $i=0,1,\ldots,N-1$, and $R_{i}=t_{i}/\sigma_{t_{i}}$. Collecting everything above, an equivalent reformulation of (19) for RF sampling is
@@
where $\eta_{i}=1-\frac{R_{i}^{2}}{R_{i+1}^{2}},\psi_{i}=\frac{R_{i}^{2}}{R_{i+1}^{2}}.\frac{1-R_{i+1}^{2}}{1-R_{i}^{2}}.\left({1-\frac{R_{i}^{2}}{R_{i+1}^{2}}}\right)$, and $\widehat{s}_{t_{i}}(x)=\frac{t_{i}\widehat{v}_{t_{i}}(x)-x}{1-t_{i}}$. See Appendix B.3 for details. Lastly, we impose an assumption on the approximation error of $\widehat{s}_{t_{i}}$.
@@
Define $(\varepsilon^{\rm s}_{i})^{2}:=\mathbb{E}\|s_{t_{i}}(X_{t_{i}})-\widehat{s}_{t_{i}}(X_{t_{i}})\|_{2}^{2}$. Then, we assume that $\frac{1}{N}\sum_{i=1}^{N}(\varepsilon^{\rm s}_{i})^{2}\leq\varepsilon_{\rm s}^{2}.$
Using the equivalence between (19) and (21), we obtain convergence guarantee for Stoc-RF, which the second main result of this paper.
@@
Note that $t_{N-1}=\frac{\sqrt{\omega_{1}}}{\sqrt{\omega_{1}}+\sqrt{1-\omega_{1}}}\approx 1$. Therefore, $\text{Law}(X_{t_{N-1}})\approx p_{1}$. In comparison with the results in Theorem 4.5, Stoc-RF sampler does not require any conditions on higher-order derivatives of the score or velocity field, as in Assumption 4.4. That is, Stoc-RF sampler can achieve similar performance under less stringent requirements. This is primarily due to the Gaussian perturbation term in (21) which acts as a corrector step. The proof is deferred to Appendix B.3.
@@
Using the equivalence relation (14), one can also construct a deterministic RF sampler using the DDIM sampler considered in Liang et al. (2025, Theorem 1) and the same time change transformation considered in this section. The time schedule in this case also follows a U-shaped distribution, however, it is different from (15) (See Appendix B.4). One can show that the new sampler also adapts to the low-dimensionality of $p_{1}$. However, we omit this discussion for brevity as the RF sampler in Theorem 4.5 already adapts to low-dimension with simpler time-steps.
Figure 3: (a)-(b) TV  distance between blurred low-rank Gaussian distribution and the generated samples by the RF sampler under uniform and non-uniform time-grid with varying $d$. (c)-(d) TV  distance between blurred low-rank Gaussian distribution and the generated samples by the RF sampler under uniform and non-uniform time-grid with varying $d$.

### v2

@@ 4.1 Low-dimensionality of target distribution
To formalize the notion of low-dimensionality of $p_{1}$, we turn to the geometric complexity of its support $\mathcal{X}$ as measured by its metric entropy.
@@
Let $\epsilon_{0}=N^{-c_{\epsilon_{0}}}$ for some sufficiently large $c_{\epsilon_{0}}>0$. The $\epsilon_{0}$ covering number of $\mathcal{X}$ with respect to the Euclidean norm, $\mathcal{N}(\mathcal{X},\|\cdot\|_{2},\epsilon_{0})$, satisfies $\log\mathcal{N}(\mathcal{X},\|\cdot\|_{2},\epsilon_{0})\leq C_{\rm cover}k\log N,$ for some constant $C_{\rm cover}>0$ and integer $k$. The quantity $k$ is referred as the intrinsic-dimension of $\mathcal{X}$.
@@
There exists a universal constant $c_{R}>0$ such that $\sup_{x\in\mathcal{X}}\|x\|_{2}\leq R,\text{where $R=N^{c_{R}}$}.$
The bounded support assumption is not as stringent as it may appear at first, as the diameter of $\mathcal{X}$ is allowed to scale polynomially (with arbitrarily large degree) in the number of iterations of the sampler, which accommodates a wide range of practical applications, such as image generation.
@@ 4.2 Assumptions on velocity approximation error
Our next set of assumptions are concerned with the approximation error of the estimated velocity field $\hat{v}_{t}$.
@@
Define $(\varepsilon^{\rm v}_{i})^{2}:=\mathbb{E}\|v_{t_{i}}(X_{t_{i}})-\widehat{v}_{t_{i}}(X_{t_{i}})\|_{2}^{2}$. Then, $\frac{1}{N}\sum_{i=1}^{N}(\varepsilon^{\rm v}_{i})^{2}\leq\varepsilon_{\rm v}^{2}.$
The above assumption essentially requires a good control on the mean-squared error of the estimated velocity field. Similar assumptions have been considered in Bansal et al. (2025); Guan et al. (2025) to prove Wasserstein convergence of RF. In the context of score-matching algorithms, similar assumptions related to the score-approximation error have also appeared in prior works such as Benton et al. (2023); Chen et al. (2023); Li et al. (2024b); Liang et al. (2025).
@@
    Let $(\varepsilon^{\rm J,1}_{i})^{2}:=\mathbb{E}\|\nabla v_{t_{i}}(X_{t_{i}})-\nabla\widehat{v}_{t_{i}}(X_{t_{i}})\|_{F}^{2}$. Then, $N^{-1}\sum_{i=1}^{N}(\varepsilon^{\rm J,1}_{i})^{2}\leq\varepsilon_{\rm J,1}^{2}$. Also, we assume $\sup_{x\in\mathbb{R}^{d}}\eta_{i}\|\nabla v_{t_{i}}(x)-\nabla\widehat{v}_{t_{i}}(x)\|_{{\rm op}}\leq 1/8$ for all $0\leq i\leq N-2$.
@@
Nearly identical assumptions have been used in Liang et al. (2025) in the context of DDIM. Intuitively, an adequate control of the higher-order approximation terms mitigates the propagation of error during sampling, as the deterministic samplers are unable to self-correct the path of the flow.
@@ 4.3 Convergence rate of rectified flow
Our convergence guarantees for the RF sampler (5) depend crucially on a carefully constructed U-shaped time schedule $\{t_{i}\}_{i=0}^{N}$, which we describe next. For given, user-specified values $h\in(0,1)$ and $\delta\geq 0$, we consider the non-uniform time partition
@@
We remark that the proposed time schedule is practicable, as it does not depend on any properties of $p_{1}$. Typically, $\delta$ is chosen to be small, as this give rise to a U-shaped time discretization as shown in Figure 2. This type of choice is not unusual (Lee et al., 2024), as it yields more accurate estimates $\widehat{v}_{t}$ during training when $t\approx 0$ and $t\approx 1$. Prior empirical work by non-uniform time discretizations have also been considered by Huang et al. (2024); Li and Yan (2024); Liang et al. (2025) in the context DDPM and DDIM to establish low-dimensional adaptation of the models. It is important to emphasize that the specific form of the time schedule (15), while conforming to best empirical practices, is theoretically grounded, and it was derived directly as a result of our technical analysis of the RF dynamics (see Step 7 of the proof of Theorem 4.5, Appendix B.2). We now present the main convergence result for the RF sampler.
@@
The dependence on $\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}$. The upper bound in Theorem 4.5 contains a term involving $\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}$, which even for data supported on manifolds could scale linearly in the ambient dimension $d$. However, in practice one can choose the free parameter $\delta$ to be e.g. $O(d^{-1})$, so that the term becomes negligible. With this choice of $\delta$, last term in (17) scales as $d^{2}N^{-10}$ which is also negligible in practice.
@@
Higher-order approximation error terms. The upper bound in Theorem 4.5 is heavily dependent on the higher-order approximation errors of $\widehat{v}_{t}$. These terms are required to be small as deterministic samplers are unable to self-correct their trajectories. In addition, at $t=0$, the true drift $v_{0}(x)=\mu_{1}-x$ can be estimated by $\widehat{v}_{0}(x)=n^{-1}\sum_{i\in[n]}{X_{1}^{(i)}}-x$, where $\{X_{1}^{(i)}\}_{i\in[n]}\overset{i.i.d.}{\sim}p_{1}$. With this choice of $\widehat{v}_{0}$, we have $\varepsilon^{\rm J,1}_{0}=\varepsilon^{\rm J,2}_{0}=\varepsilon^{\rm H}_{0}=0$, and $\varepsilon^{\rm v}_{0}=\sqrt{\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}/n}$.
@@ 4.4 Convergence rate of STOC-RF
In this section, we will leverage the equivalence between DDPM and STOC-RF to design an SDE-based sampler that also adapts to the low-dimensionality of $p_{1}$. We begin with the solution $\left\{{Y^{\prime}}\right\}_{\tau\geq 0}$ to the forward process (11). We also consider the time-change map $\tau\in[0,\infty)\mapsto t(\tau)=\frac{\sqrt{\omega_{\tau}}}{\sqrt{\omega_{\tau}}+\sqrt{1-\omega_{\tau}}}.$ In light of (14) and Proposition 3.5, we note that $\left\{{Y^{\prime}_{\tau}}\right\}_{\tau\geq 0}$ and $\{\tilde{Z}_{t}\}_{t\in[0,1]}$ (recall 6) are equivalent under the time change $\tau\mapsto t(\tau)$, and we have $\frac{Y^{\prime}_{\tau}}{\sqrt{\omega_{\tau}}}\overset{d}{=}\frac{\tilde{Z}_{t(\tau)}}{t(\tau)}$. In practice, we simulate the forward process by discretizing $\tau$ to positive integers and setting specific values for $\omega_{\tau}$. Here, we let $\omega_{\tau}=\prod_{j=1}^{\tau}\alpha_{j}$ with $\alpha_{j}$’s defined implicitly as
$\beta_{1}=1-\alpha_{1}=1/N^{c_{0}},$ (18)
@@
for a user-specified choice of the parameters $c_{0},c_{1}>0$. The time schedule above is very similar to the ones considered by Potaptchik et al. (2024); Li et al. (2024b); Li and Yan (2024); Huang et al. (2024); Liang et al. (2025) to analyze the convergence properties of DDPM. Next, we recall the DDPM sampler considered in Liang et al. (2025); Huang et al. (2024); Li and Yan (2024):
@@
$\delta_{\tau}=1-\alpha_{\tau},\nu_{\tau}=\sqrt{\frac{(\alpha_{\tau}-\omega_{\tau})(1-\alpha_{\tau})}{1-\omega_{\tau}}},$
@@
where $\widehat{s}_{Y^{\prime}_{\tau}}$ is an approximate score of $Y^{\prime}_{\tau}$. To provide some intuition behind this sampler, we recall that $1-\alpha_{\tau}=\beta_{\tau}\approx 0$. Hence, $1/\sqrt{\alpha_{\tau}}\approx 1+\beta_{\tau}/2,\delta_{\tau}/\sqrt{\alpha_{\tau}}\approx\beta_{\tau}/2$, and $\nu_{\tau}/\sqrt{\alpha_{\tau}}\approx\sqrt{\beta_{\tau}}$. Substituting this in (19), we get
$\hat{Y}^{\prime}_{\tau-1}\approx\hat{Y}^{\prime}_{\tau}+\frac{\beta_{\tau}}{2}\left\{{\hat{Y}^{\prime}_{\tau}+2\widehat{s}_{Y^{\prime}_{\tau}}(\hat{Y}^{\prime}_{\tau})}\right\}+\sqrt{2\left({\frac{\beta_{\tau}}{2}}\right)}\xi_{\tau},$ (20)
@@
Therefore, we have $Y^{\prime}_{\tau}\overset{d}{=}\tilde{Z}_{t(\tau)}/\sigma_{t(\tau)}$, and $s_{Y^{\prime}_{\tau}}(y)=\sigma_{t(\tau)}s_{t(\tau)}(\sigma_{t(\tau)}y)$. We also define $t_{i}:=t(N-i)$ for $i=0,1,\ldots,N-1$, and $R_{i}=t_{i}/\sigma_{t_{i}}$. Collecting everything above, we arrive at an equivalent reformulation of (19) for RF sampling of the form
@@
where $\eta_{i}=1-\frac{R_{i}^{2}}{R_{i+1}^{2}},\psi_{i}=\frac{R_{i}^{2}}{R_{i+1}^{2}}\frac{1-R_{i+1}^{2}}{1-R_{i}^{2}}\left({1-\frac{R_{i}^{2}}{R_{i+1}^{2}}}\right)$, and $\widehat{s}_{t_{i}}(x)=\frac{t_{i}\widehat{v}_{t_{i}}(x)-x}{1-t_{i}}$. See Appendix B.3 for details. Lastly, we impose an assumption on the approximation error of $\widehat{s}_{t_{i}}$.
@@
Define $(\varepsilon^{\rm s}_{i})^{2}:=\mathbb{E}\|s_{t_{i}}(X_{t_{i}})-\widehat{s}_{t_{i}}(X_{t_{i}})\|_{2}^{2}$. Then, $\frac{1}{N}\sum_{i=1}^{N}(\varepsilon^{\rm s}_{i})^{2}\leq\varepsilon_{\rm s}^{2}.$
Leveraging the equivalence between (19) and (21), we obtain convergence guarantee for STOC-RF, which the second main result of this paper.
@@
Note that $t_{N-1}=\frac{\sqrt{\omega_{1}}}{\sqrt{\omega_{1}}+\sqrt{1-\omega_{1}}}\approx 1$. Therefore, $\text{Law}(X_{t_{N-1}})\approx p_{1}$. In comparison with Theorem 4.5, STOC-RF sampler does not require any conditions on higher-order derivatives (4.4) of the score or velocity field. That is, STOC-RF sampler can achieve similar performance under less stringent requirements. This is primarily due to the Gaussian perturbation term in (21) which acts as a corrector step. The proof is in Appendix B.3.
@@
Using the equivalence relation (14), it is also possible to directly relate the deterministic RF sampler to the DDIM sampler studied by Liang et al. (2025, Theorem 1) (Appendix B.4) via the same time change transformation considered in this section. The resulting time schedule also follows a U-shaped distribution and the sampler will also adapts to the low-dimensionality of $p_{1}$.